OpenStudy (saifoo.khan):

The sum of five integers is 190. If one of these numbers is removed, the average of the four remaining numbers is between 34.5 and 36, inclusive. What is one possible value for the number that was removed?

OpenStudy (anonymous):

Let the five numbers be a,b,c,d,e.And let e be the number that is removed. a+b+c+d+e = 190. \[34.5\le (190-e)/4 \le 36 \] On simplifying the above inequality,we get \[138 \le 190-e \le 144\] \[-52 \le -e \le -46\] \[46 \le e \le 52\] So, the value of the missing integer e can be between 46 and 52 inclusive.

OpenStudy (anonymous):

I forgot to mention this: The average after removing e is (190-e)/4 .

OpenStudy (saifoo.khan):

Wow, thanks a ton! @kartiksriramk

OpenStudy (anonymous):

You're welcome man. Nice question by the way.